# UniChest: Conquer-and-Divide Pre-training for Multi-Source Chest X-Ray Classification

### 18 Dec 2023 v1

@@ 3.3.2 The Training at the “Conquer" Stage
where $\bar{\boldsymbol{I}}_{i}$ is the average pooling of $\boldsymbol{I}_{i}$ along the first two dimensions so that we reduce it to the same dimension of $\boldsymbol{T}_{i}$, and $\tau$ is temperature with the default setting 1.0 following (Chen et al., 2020). Regarding $\mathcal{L}_{\text{BCE}}$, we only enforce the supervision on the first query network and compute the multi-label cross-entropy loss at this stage. Note that, samples in some domains may have smaller label space than the union label space of all sources. In this case, we neglect the computation between prediction and missing classes of these samples, even if it may belong to them (but are unobserved). Finally, we sum up $\mathcal{L}_{\text{BCL}}$ and $\mathcal{L}_{\text{BCE}}$ (only involves the first query network) as the overall loss for each mini-batch at the “Conquer" stage as follows
@@
3.3.3 The Training at the “Divide" Stage
Previous stage follows the classical pre-training spirit, where we treat multiple sources equally, and expect that the model can sufficiently learn the multi-source common patterns. At the “Divide" phase, we mainly explore to mediate the negative impact induced by the source heterogeneity. Concretely, on the basis of the warming-up during the “Conquer" stage, we train all the query networks with freezing all the other parts of the model. Here, the multi-label cross-entropy loss will be implemented on the ensemble prediction $s^{\text{All}}$ from $\{s^{k}\}_{k=1}^{K}$, unlike using $s^{1}$ in Eq. (4). The final prediction is characterized by the following equation
where $\lambda$ is the hyperparameter to balance the training of the first query network and the remaining query networks, and $\omega_{i}^{k}$ is the learnable weight to summarize the contribution of the source-specific modules for the diagnosis prediction. Note that, the intuition behind Eq. (5) is to inherit the training gains at the “Conquer" stage by $\lambda$, and simultaneously squeeze the source-specific patterns to the other query networks by $\omega_{i}^{k}$.
Source Contrastive Learning. In Eq. (5), we introduce the learnable weights to incorporate the source-specific patterns into the other query networks. However, without any guidance, it is challenging to achieve this goal by optimization. Here, we introduce a source contrastive learning to promote the desire. Specifically, we transform the global visual representation $\bar{\boldsymbol{I}}_{i}$ into a $K-1$ simplex via one MLP layer $\Phi_{1}(\cdot)$ as the weight vector, namely, $\boldsymbol{\omega}_{i}=\Phi_{1}(\bar{\boldsymbol{I}}_{i})\in\mathbb{R}^{K-1}$. Then, $\boldsymbol{\omega}_{i}$ is projected into a higher dimension space by another MLP layer $\Phi_{2}(\cdot)$, denoted as $\bar{\boldsymbol{\omega}}_{i}=\Phi_{2}(\boldsymbol{\omega}_{i})$, to perform a source contrastive learning with the guidance of source id in the following
Above contrastive learning helps us to learn the similar weight vector for the samples from the same source, which makes the source-specific patterns learned by the similar query networks. Under this mechanism, we naturally learn an automatic optimal assignment for the remaining query networks to overcome the source heterogeneity issue. Then, with Eq. (5) and Eq. (6), the overall loss at the “Divide" stage can be formulated as
Table 1: Statistics of datasets for pre-training and downstream.
Pre-training MIMIC-CXR 348900 41 Northeast USA MIT 2011-2016
ChestX-ray14 98637 14 Northeast USA NIH 1992-2015
CheXpert 223414 14 Western USA Stanford Hospital 2002-2017
VinDr-CXR 15000 28 Vietnam VinBrain 2018-2020
Multi-CXR 685951 60 USA & Vietnam MIT, etc. 1992-2020
Downstream Shenzhen 662 1 Southeast China Guangdong Medical College Sep. 2012
Open-I 3547 14 USA NLM /
SIIM-ACR 2135 1 Northeast USA Kaggle /
PadChest 39053 193 Spain University of Alicante 2009-2017
ChestX-Det10 542 / Northeast USA Deepwise 1992-2015
@@ 3.3.4 Difference from the pre-training with fine-tuning
The proposed Conquer-and-Divide pre-training is intrinsically different from the ordinary fine-tuning after pre-training, although both of them are a two-stage process. First, the “Divide" training of UniChest is towards all multi-source data, while the fine-tuning after pre-training is narrowed down to a single-source scenario. This leads to the distinction in the generalization ability of two frameworks, where UniChest is significantly stronger than pre-training with fine-tuning as supported by Table 2. Second, the “Divide" stage is to properly mediate the heterogeneity patterns in a broad sense, which does not mean these patterns are not related to the generalization. Conversely, the downstream task can still apply the ordinary fine-tuning ways to leverage these patterns by automatic activated weights for diagnosis, and achieve the fast adaptation. Finally, when it comes to the experimental comparison, we mainly focus on the zero-shot performance of pre-training, instead of the performance by the fine-tuning, which follows the prevalent pre-training spirit that tries to avoid the expensive tuning cost for downstream tasks as much as possible.
@@ 4.1 Datasets
We combine some common CXR datasets containing MIMIC-CXR (Johnson et al., 2019), ChestX-ray14 (Wang et al., 2017), CheXpert (Irvin et al., 2019) and VinDr-CXR (Nguyen et al., 2022) for pre-training, namely Multi-CXR, which yields a grand total of 685,951 Chest X-Ray images.
For evaluating the generalization performance of the pre-trained model, we adopt Shenzhen (Jaeger et al., 2014), Open-I (Demner-Fushman et al., 2016), SIIM-ACR Pneumothorax (kag, ) and PadChest (Demner-Fushman et al., 2016) for zero-shot classification and ChestX-Det10 (Liu et al., 2020) which is the subset of ChestX-ray14 for intuitive lesion grounding. More detailed statistics of pre-training and downstream datasets can be found in Table 1. We can observe that significant diversity exists in the collection time, population distribution, disease category coverage and sample scales among different datasets.
@@ 4.2 Implementation Details
The pre-training process of UniChest is conducted on a single NVIDIA A100 GPU for 30 epochs in the first “Conquer” stage and 20 epochs in the second “Divide” stage. The starting checkpoint for the second stage is the one that performed the best on the ChestX-ray14 validation set. The depth of each query network in MoE-QN Module is set to 4. The hyperparameter $\lambda$ which weights the first query network during the “Divide” stage is set to be $0.5$ and the total number of query network $K$ is $4$ in default. We set the training batch size as 32 and resize input images as 512 $\times$ 512. We adopt the AdamW optimizer in conjunction with the cosine annealing scheduler for managing the learning rate, where the initial learning rate is $1\times 10^{-5}$. To perform data augmentation, we additionally utilize the Fourier amplitude mixup method, which has been demonstrated to be effective for medical imaging data (Zhang et al., 2022a; Liu et al., 2021). Besides, we utilize the ASL loss (Ridnik et al., 2021) as $\mathcal{L}_{\text{BCE}}$. These techniques have demonstrated promising outcomes in the analysis of CXR data.
@@
4.3 Evaluation Metrics
Some common metrics for multi-label classification are adopted to evaluate the model performance, i.e., area under curve (AUC), F1 score, accuracy (ACC) and average precision (AP) for each category and their average value, namely, average AUC (aAUC), average F1 (aF1), average accuracy (aACC) and mean average precision (mAP), for comprehensive comparison. Following the strategy of MedKLIP (Wu et al., 2023), the final binary threshold of each category prediction is the value when the maximum F1 score is achieved and ACC metric also adopts this threshold.
4.4 Baselines
We consider a wide range of baselines in Chest X-Ray pre-training, including ConVIRT (Zhang et al., 2022b), GLoRIA (Huang et al., 2021), BioVIL (Boecking et al., 2022), CheXzero (Tiu et al., 2022), MedKLIP (Wu et al., 2023), and KAD (Zhang et al., 2023).
ConVIRT trains two modality-specific encoders by bidirectional contrastive loss to learn visual representations. GLoRIA utilizes both global and fine-grained features for medical VLP. BioVIL proposes a radiology-specific text encoder for the subsequent classical pipeline of VLP. CheXzero retrains one CLIP model with a corpus of the medical domain. MedKLIP designs one novel entity extraction and transition module to inject domain-specific knowledge into the process of VLP. KAD incorporates a medical knowledge graph to further improve the capability of current VLP models in Chest X-Ray, showing SOTA performance on some common public benchmarks.
@@ 4.5.1 In-domain evaluation on pre-training datasets
We compare UniChest with other baselines on the test set of pre-training datasets in Table 3. UniChest demonstrates a notable improvement in various evaluation metrics compared to the best baseline models. Specifically, for ChestX-ray14, UniChest shows a significant improvement of 6.51%, 9.30%, 3.60% and 10.51% respectively as to aAUC, aACC, aF1 and mAP than the best baseline. In terms of fine-grained classification for each category, UniChest demonstrates persistent gains as shown in Fig, 4 (left). For CheXpert, UniChest outperforms 6.16%, 9.79%, 2.40% and 11.28%. In the case of VinDr-CXR, UniChest surpasses the best baselines by 12.08%, 10.68%, 7.77% and 12.05%. By introducing multi-source CXR datasets and the Conquer-and-Divide pre-training framework, the diagnosis ability of VLP model has been significantly enhanced.
@@ 4.5.3 Qualitative grounding visualization
(a) aAUC Score
(b) aF1 Score
(c) aACC Score
(d) mAP Score
@@ 5.1.1 On “Conquer" stage
5.1.2 On “Divide" stage
A comparison of the second and fourth rows of Table 5 reveals that UniChest significantly outperforms the “Conquer” stage in terms of zero-shot generalization by 4.86% (AUC), 3.48% (F1), 3.19% (ACC), and 2.85% (AP) for Shenzhen, and 2.41% (AUC), 4.78% (F1), 3.13% (ACC), and 3.28% (AP) for SIIM-ACR Pneumothorax. To further present the effectiveness of the soft-gating mechanism and $\mathcal{L}_{\text{SCL}}$ in the MoE-QN structure during the “Divide” stage, we present the results of removing the weight generation process and replacing it with equal weights in the third row of Table 5. As shown on SIIM-ACR Pneumothorax, our mechanism significantly enhances zero-shot generalization.

### 22 Mar 2024 v2

@@ 3.3.2 The Training at the “Conquer” Stage
where $\bar{\boldsymbol{I}}_{i}$ is the average pooling of $\boldsymbol{I}_{i}$ along the first two dimensions so that we reduce it to the same dimension of $\boldsymbol{T}_{i}$, and $\tau$ is temperature with the default setting 1.0 following [40]. Regarding $\mathcal{L}_{\text{BCE}}$, we only enforce the supervision on the first query network and compute the enhanced multi-label cross-entropy loss at this stage. Note that, samples in some domains may have smaller label space than the union label space of all sources. In this case, we neglect the computation between prediction and missing classes of these samples, even if it may belong to them (but are unobserved). Finally, we sum up $\mathcal{L}_{\text{BCL}}$ and $\mathcal{L}_{\text{BCE}}$ (only involves the first query network) as the overall loss for each mini-batch at the “Conquer” stage as follows
@@
Table 2: Comparison with other state-of-the-art methods on fine-tuning classification task with different training data portions on SIIM-ACR Pneumothorax. FT denotes fine-tuning pre-trained model, ZS denotes zero-shot classification.
Data Portion 1% 10% 100%
[…]
ConVIRT (FT) 0.7134 0.4651 0.6375 0.4224 0.7826 0.5346 0.7190 0.4640 0.8989 0.6858 0.8351 0.7039
GLoRIA (FT) 0.7439 0.4948 0.7030 0.4595 0.8538 0.6091 0.7874 0.6250 0.9018 0.6898 0.8450 0.7269
BioVIL (FT) 0.6947 0.4685 0.6244 0.3816 0.7775 0.5393 0.7372 0.5183 0.8688 0.6364 0.8004 0.6624
MedKLIP (FT) 0.8527 0.5952 0.7925 0.5858 0.9071 0.6845 0.8389 0.6823 0.9194 0.7246 0.8689 0.7510
KAD (FT) 0.8541 0.5810 0.8097 0.4822 0.9023 0.6498 0.8490 0.6506 0.9203 0.6788 0.8655 0.7079
UniChest (ZS) 0.9260 0.6928 0.8807 0.6929 0.9260 0.6928 0.8807 0.6929 0.9260 0.6928 0.8807 0.6929
UniChest (FT) 0.9375 0.7182 0.8801 0.7326 0.9418 0.7314 0.8881 0.7508 0.9437 0.7254 0.8885 0.7696
3.3.3 The Training at the “Divide” Stage
Previous stage follows the classical pre-training spirit, where we treat multiple sources equally, and expect that the model can sufficiently learn the multi-source common patterns. At the “Divide” phase, we mainly explore to mediate the negative impact induced by the source heterogeneity. Concretely, on the basis of the warming-up during the “Conquer” stage, we train all the query networks (MoE-QN Module) with freezing all the other parts of the model. In the MoE-QN Module, the new query networks introduced at the this stage are initialized from randomness. Here, the multi-label cross-entropy loss will be implemented on the ensemble prediction $s^{\text{All}}$ from $\{s^{k}\}_{k=1}^{K}$, unlike using $s^{1}$ in Eq. (4). The final prediction is characterized by the following equation
where $\lambda$ is the hyperparameter to balance the training of the first query network and the remaining query networks, and $\omega_{i}^{k}$ is the learnable weight to summarize the contribution of the source-specific modules for the diagnosis prediction. Note that, the intuition behind Eq. (5) is to inherit the training gains at the “Conquer” stage by $\lambda$, and simultaneously squeeze the source-specific patterns to the other query networks by $\omega_{i}^{k}$.
Source Contrastive Learning. In Eq. (5), we introduce the learnable weights to incorporate the source-specific patterns into the newly introduced query networks during the “Divide” Stage. However, without any guidance, it is challenging to achieve this goal by optimization. Here, we introduce a source contrastive learning to promote the desire. Specifically, we transform the global visual representation $\bar{\boldsymbol{I}}_{i}$ into a $K-1$ simplex via one-layer MLP $\Phi_{1}(\cdot)$ and Softmax layer as the weight vector, namely, $\boldsymbol{\omega}_{i}=\Phi_{1}(\bar{\boldsymbol{I}}_{i})\in\mathbb{R}^{K-1}$. Then, $\boldsymbol{\omega}_{i}$ is projected into a higher dimension space by another one-layer MLP $\Phi_{2}(\cdot)$, denoted as $\bar{\boldsymbol{\omega}}_{i}=\Phi_{2}(\boldsymbol{\omega}_{i})$, which is the high-dim vector to perform a source contrastive learning with the guidance of source id $l$ in the following
In Eq. (6), the numerator ${\sum_{l_{j}=l_{i}}^{j\neq i}{\rm e}^{\bar{\boldsymbol{\omega}}_{i}^{\top}\bar{\boldsymbol{\omega}}_{j}/\tau}}$ computes the sum of exponential similarities of the high-dim weight vectors $\bar{\boldsymbol{\omega}}$ between sample $i$ and all other samples $j$ from the same source $l$, encouraging to pull the samples with the same source identify closer together. The denominator ${\sum_{m\neq i}^{M}{\rm e}^{\bar{\boldsymbol{\omega}}_{i}^{\top}\bar{\boldsymbol{\omega}}_{m}/\tau}}$ calculates the sum of exponential similarities of the high-dim weight vectors between sample $i$ and all other samples in the batch, providing a normalization factor to ensure stability of the loss. Generally, above contrastive learning helps us to learn the similar weight vector for the samples from the same source, which makes the source-specific patterns learned by the similar query networks. Under this mechanism, we naturally learn an automatic optimal assignment for the remaining query networks to overcome the source heterogeneity issue. Then, with Eq. (5) and Eq. (6), the overall loss at the “Divide” stage can be formulated as
Table 3: Statistics of datasets for pre-training and downstream.
Pre-training MIMIC-CXR [8] 348900 41 Northeast USA MIT 2011-2016
ChestX-ray14 [12] 98637 14 Northeast USA NIH 1992-2015
CheXpert [13] 223414 14 Western USA Stanford Hospital 2002-2017
VinDr-CXR [14] 15000 28 Vietnam VinBrain 2018-2020
Multi-CXR 685951 60 USA & Vietnam MIT, etc. 1992-2020
Downstream Shenzhen [15] 662 1 Southeast China Guangdong Medical College Sep. 2012
Open-I [16, 41] 3547 14 USA NLM /
SIIM-ACR [17, 6] 2135 1 Northeast USA Kaggle /
PadChest [16] 39053 193 Spain University of Alicante 2009-2017
ChestX-Det10 [42] 542 / Northeast USA Deepwise 1992-2015
@@ 3.3.4 Difference from the pre-training with fine-tuning
The proposed Conquer-and-Divide pre-training is intrinsically different from the ordinary fine-tuning after pre-training, although both of them are a two-stage process. First, the two-stage training of UniChest is towards all multi-source data, while the fine-tuning after pre-training is narrowed down to a single-source scenario. When the target dataset’s corresponding training set is seen during multi-source pre-training, this leads to the distinction in the generalization ability of two frameworks, where UniChest is significantly stronger than pre-training with fine-tuning as supported by Table 1. Second, the “Divide” stage is to properly mediate the heterogeneity patterns in a broad sense, which does not mean multi-source pre-training contradicts with the conventional pre-training and fine-tuning diagram. Conversely, the downstream task can still apply the ordinary fine-tuning ways to achieve fast adaptation for unseen sources. For instance, we fine-tune the pre-trained UniChest on SIIM-ACR Pneumothorax [17] with different training data ratios following MedKLIP [6]. Overall, UniChest’s zero-shot capability is already comparable to the 100% fine-tuned performance of other models and the diagnostic performance is further enhanced after further fine-tuning as shown in Table 2. Finally, when it comes to the subsequent experimental comparison, we mainly focus on the zero-shot performance of pre-training, instead of the performance by the fine-tuning, which follows the prevalent pre-training spirit that tries to avoid the expensive tuning cost for downstream tasks as much as possible.
@@ 4.1 Datasets
We combine some common CXR datasets containing MIMIC-CXR [8], ChestX-ray14 [12], CheXpert [13] and VinDr-CXR [14] for pre-training, namely Multi-CXR, which yields a grand total of 685,951 Chest X-Ray images.
For evaluating the generalization performance of the pre-trained model, we adopt Shenzhen [15], Open-I [16], SIIM-ACR Pneumothorax [17] and PadChest [16] for zero-shot classification and ChestX-Det10 [42] which is the subset of ChestX-ray14 for intuitive lesion grounding. More detailed statistics of pre-training and downstream datasets can be found in Table 3. We can observe that significant diversity exists in the collection time, population distribution, disease category coverage and sample scales among different datasets.
@@ 4.2 Implementation Details
The pre-training process of UniChest is conducted on a single NVIDIA A100 GPU for 30 epochs in the first “Conquer” stage and 20 epochs in the second “Divide” stage. The starting checkpoint for the second stage is the one that performed the best on the ChestX-ray14 validation set. The number of transformer decoder layers of each query network in MoE-QN Module is set to 4. The hyperparameter $\lambda$ which weights the first query network during the “Divide” stage is set to be $0.5$ and the total number of query network $K$ is $4$ in default. The dimension of $\bar{\boldsymbol{\omega}}$ which is the projection of the MLP layer $\Phi_{2}(\cdot)$ is set to be 32 during source contrastive learning. The temperature parameter $\tau$ in Eq. (6) is set to 1.0. We set the training batch size as 32 and resize input images as 512 $\times$ 512. We adopt the AdamW optimizer in conjunction with the cosine annealing scheduler for managing the learning rate, where the initial learning rate is $1\times 10^{-5}$. To perform data augmentation, we additionally utilize the Fourier amplitude mixup method, which has been demonstrated to be effective for medical imaging data [43, 44]. Besides, we utilize the ASL loss [45] as $\mathcal{L}_{\text{BCE}}$ to promote balanced training. These techniques have demonstrated promising outcomes in the analysis of CXR data. For MedKLIP-multi and KAD-multi, we strictly follow the training hyper-parameters and implementation details described in their original papers and official codes. Their pre-training dataset is also Multi-CXR that is identical to UniChest. In MedKLIP-multi, for samples without location annotations (excluding MIMIC-CXR), we mask the calculation of location contrastive loss directly.
4.3 Evaluation Metrics
Some common metrics for multi-label classification are adopted to evaluate the model performance, i.e., area under curve (AUC), F1 score, accuracy (ACC) and average precision (AP) for each category and their average value, namely, average AUC (aAUC), average F1 (aF1), average accuracy (aACC) and mean average precision (mAP), for comprehensive comparison. Following the strategy of MedKLIP [6], the final binary threshold of each category prediction is the value when the maximum F1 score is achieved and ACC metric also adopts this threshold.
4.4 Baselines
We consider a wide range of baselines for CXR pre-training, including ConVIRT [33], GLoRIA [34], MedCLIP [35], BioVIL [4], CheXzero [5], MedKLIP [6], and KAD [7].
ConVIRT trains two modality-specific encoders by bidirectional contrastive loss to learn visual representations. GLoRIA utilizes both global and fine-grained features for medical VLP. MedCLIP [35] trains one CLIP-based framework on two CXR datasets including MIMIC-CXR and CheXpert. BioVIL proposes a radiology-specific text encoder for the subsequent classical pipeline of VLP. CheXzero retrains one CLIP model with a corpus of the medical domain. MedKLIP designs one novel entity extraction and transition module to inject domain-specific knowledge into the process of VLP. KAD incorporates a medical knowledge graph to further improve the capability of current VLP models in Chest X-Ray, showing SOTA performance on some common public benchmarks.
@@ 4.5.1 In-domain evaluation on pre-training datasets
We compare UniChest with other baselines on the test set of pre-training datasets in Table 4. UniChest demonstrates a notable improvement in various evaluation metrics compared to the best baseline models. Specifically, for ChestX-ray14, UniChest shows a significant improvement of 6.51%, 9.30%, 3.60% and 10.51% respectively as to aAUC, aACC, aF1 and mAP than the best baseline. In terms of fine-grained classification for each category, UniChest demonstrates persistent gains as shown in Fig. 4 (left). For CheXpert, UniChest outperforms 6.16%, 9.79%, 2.40% and 11.28%. In the case of VinDr-CXR, UniChest surpasses the best baselines by 12.08%, 10.68%, 7.77% and 12.05%. By introducing multi-source CXR datasets and the Conquer-and-Divide pre-training framework, the diagnosis ability of VLP model has been significantly enhanced.
@@ 5.1.1 On “Conquer” stage
5.1.2 On “Divide” stage
A comparison of the second and fourth rows of Table 6 reveals that UniChest significantly outperforms the “Conquer” stage in terms of zero-shot generalization by 4.86% (AUC), 3.48% (F1), 3.19% (ACC), and 2.85% (AP) for Shenzhen, and 2.41% (AUC), 4.78% (F1), 3.13% (ACC), and 3.28% (AP) for SIIM-ACR Pneumothorax. To further present the effectiveness of the soft-gating mechanism and $\mathcal{L}_{\text{SCL}}$ in the MoE-QN structure during the “Divide” stage, we present the results of removing the weight generation process and replacing it with equal weights in the third row of Table 6. As shown on SIIM-ACR Pneumothorax, our mechanism significantly enhances zero-shot generalization. In addition to the performance improvement, it’s also important to note that the introduction of the MoE-QN structure at the “Divide” stage does not significantly increase the computational cost compared to previous single-stage pre-training frameworks. Concretely, the increase in computational complexity is marginal, growing only from 44.3 GFLOPs in KAD to 45.5 GFLOPs in UniChest, which constitutes an increase of approximately 2.7%.
(a) aAUC Score
(b) aF1 Score
(c) aACC Score
(d) mAP Score
Table 7: Ablation study on the role of the report. Report in the first row means adopting report entities as the input of the text encoder if the sample is with report. “Label Concat.” denotes we concatenate corresponding positive labels as the input. “Partial” means adopting “Label Concat.” only for samples without reports while “All” means adopting “Label Concat.” regardless of whether accompanied by reports or not.
Report Label Concat. aAUC aF1 aACC mAP
Partial All
0.8716 0.5450 0.8939 0.5175
0.8742 0.5488 0.8926 0.5217
0.8733 0.5624 0.8995 0.5458
0.8831 0.5687 0.9039 0.5508
@@ 5.2 The Robustness under Different Hyper-parameters
5.3 The Role of Report in Pre-training
As stated in Section 3, following the paradigm of vision-language pre-training, we use entities extracted from the report by NER tool as the input of the text encoder $\Phi_{\text{text}}$ if the corresponding report is available, while we concatenate their positive labels as input content for those without reports. In our multi-source pre-training dataset Multi-CXR, there are a total of 685,951 samples, of which 348,900 samples from MIMIC-CXR have reports, meaning that half of the samples come with reports while the rest samples have no corresponding reports. It’s worth noting that numerous datasets, which currently present only labels and images, such as CheXpert and ChestX-ray14, were originally accompanied by reports. The labels from these datasets were derived using NLP tools akin to our entity extraction method. Nevertheless, due to certain practical restrictions, including privacy concerns, these reports are not publicly available. Consequently, opting for label concatenation as a substitute is a reasonable manner. To further study the role of the report, we consider three different processing methods. First, we discard the text input directly. Second, we utilize only reports and cancel the input of positive labels for samples without reports. Third, we adopt positive label concatenation as the input for the text encoder for all samples whether they have a report or not. The overall result shown in Table 7 drops a little when we underutilize the text input, emphasizing the importance of semantically rich reports in the development of the CXR diagnosis model.